40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4997     Entry time: Wed Jul 20 10:10:19 2011
Author: Sonali 
Type: Update 
Category: Green Locking 
Subject: Weekly summary 

 I finished wih the set-up at the ETMY table. Instead of the neutral Density Filter , I put in a mirror(Y1-1037-45S)  which is reflective for IR , so that only 1% of the light is incident on the fibre  as per Rana's suggestion.

Now, the power incident on the fibre is measured to be 6 mW and the power measured out of the fibre is 2.76 mW after the necessary alignments.

On the PSL able, I have routed the beam that is coming out of the back of the PMC(instead of the dumped light from the oven to prevent any light from reflecting back into the laser), to the area where I am putting the set-up for the superposition of the PSL and the ETMX and ETMY beams.

Today I will proceed with the layout.

ELOG V3.1.3-